Solve for c.
a/b=1/c-2

Answer by checkley77(12844)

a/b=1/(c-2) a)c=b/a+2b b)c=b/a+2 c)c=2b+a/2a
a(c-2)=b
ac-2a=b
ac=2a+b
c=(2a+b)/a ans.
OR
a/b=(1/c)-2
a/b=(1-2c)/c
ac=b(1-2c)
ac=b-2bc
ac+2bc=b
c(a+2b)=b
c=b/(a+2b) ans. A).
SORRY I MESSED UP ON THE CROSS MULTIPLICATION.